OpenStudy (anonymous):

2.What is the factored form of x^2 + 6x + 8? (1 point) (x + 5)(x + 3) (x + 4)(x + 2) (x + 7)(x + 1) (x + 3)(x + 3) 3.What is the factored form of x^2 – 7x + 12? (1 point) (x – 5)(x – 3) (x – 6)(x – 1) (x – 5)(x – 2) (x – 4)(x – 3) 4.The area of a garden is given by the trinomial g^2 – 2g – 24. The garden’s length is g + 4. What is the garden’s width? g – 2 g – 6 g – 8 g + 2

OpenStudy (anonymous):

Do you know how to apporach the problem at all?

OpenStudy (anonymous):

No I don't. I haven't been to class and I don't understand any of them.

OpenStudy (anonymous):

Well, to factor a quadratic with a leading coefficient of 1 only requires you to look at the last term, and find factors of the last term which add or subtract to the ceofficient of the middle term.

OpenStudy (anonymous):

So, let's take a look at the first problem. The fzctors of the last number 8 are 1,and 8; or 2 and 4. Of those combinations, which are the ones which add to 6, the coeffiicient of your middle term?

OpenStudy (anonymous):

factors*

OpenStudy (anonymous):

2 and 4 equals 6

OpenStudy (anonymous):

That's correct. So, now we need to find the choice in your list of choices that has both a +2 and a +4 in the factors.

OpenStudy (anonymous):

Which one is it?

OpenStudy (anonymous):

So it is answer B

OpenStudy (anonymous):

You got it!. Now, as a hint to the second one, when your last number is positive but the coefficient of the middle term is negative, both of your factors must be negative.

OpenStudy (anonymous):

So you want two factors of 12 which add to give you a -7. What are they?

OpenStudy (anonymous):

#4.You have g^2-2g-24=(g+4)*width So the width has to be g+or- a number to get something starting with g^2 It has -24 so the number has to be -6 as 4*-6=-24 So the answer is g-6 You can check by expanding the bracket (g+4)(g-6)

OpenStudy (anonymous):

Thanks seiga, but I was trying to get her to understand how to factor on her own instead of just giving her the answer!

OpenStudy (anonymous):

Um, i can't find it. I'm not good with negative numbers.

OpenStudy (anonymous):

Well, can you give me two numbers which have a product of 12 and a sum of 7?

OpenStudy (anonymous):

lol sorry

OpenStudy (anonymous):

4 and 3

OpenStudy (anonymous):

Good. Now since I said they both had to be negative that means the factors of 12 that add to a negative 7 are ?

OpenStudy (anonymous):

so it would be -4 and -3 so the answer is D

OpenStudy (anonymous):

There you go!

OpenStudy (anonymous):

Thanks for helping me

OpenStudy (anonymous):

You are welcome.

OpenStudy (anonymous):

1. Which number makes this equation true? v2 + 10v + 16 = (v + 8)(v + ?) (1 point) Correct answer (1 pt) 2 Unselected answer (0 pts) 4 Unselected answer (0 pts) 6 Unselected answer (0 pts) 8 1 /1 point 2. What is the factored form of x2 + 6x + 8? (1 point) Unselected answer (0 pts) (x + 5)(x + 3) Correct answer (1 pt) (x + 4)(x + 2) Unselected answer (0 pts) (x + 7)(x + 1) Unselected answer (0 pts) (x + 3)(x + 3) 1 /1 point 3. What is the factored form of x2 – 7x + 12? (1 point) Unselected answer (0 pts) (x – 5)(x – 3) Unselected answer (0 pts) (x – 6)(x – 1) Unselected answer (0 pts) (x – 5)(x – 2) Correct answer (1 pt) (x – 4)(x – 3) 1 /1 point 4. The area of a garden is given by the trinomial g2 – 2g – 24. The garden’s length is g + 4. What is the garden’s width? (1 point) Unselected answer (0 pts) g – 2 Correct answer (1 pt) g – 6 Unselected answer (0 pts) g – 8 Unselected answer (0 pts) g + 2 1 /1 point The final score is 4/4 (100%).